### v1

# Hierarchical Diffusion Motion Planning with Task-Conditioned Uncertainty-Aware Priors

Abstract
We propose a novel hierarchical diffusion planner that embeds task and motion structure directly in the noise model. Unlike standard diffusion-based planners that use zero-mean, isotropic Gaussian noise, we employ a family of task-conditioned structured Gaussians whose means and covariances are derived from Gaussian Process Motion Planning (GPMP): sparse, task-centric key states or their associated timings (or both) are treated as noisy observations to produce a prior instance. We first generalize the standard diffusion process to biased, non-isotropic corruption with closed-form forward and posterior expressions. Building on this, our hierarchy separates prior instantiation from trajectory denoising: the upper level instantiates a task-conditioned structured Gaussian (mean and covariance), and the lower level denoises the full trajectory under that fixed prior. Experiments on Maze2D goal-reaching and KUKA block stacking show improved success rates, smoother trajectories, and stronger task alignment compared to isotropic baselines. Ablation studies indicate that explicitly structuring the corruption process offers benefits beyond simply conditioning the neural network. Overall, our method concentrates probability mass of prior near feasible, smooth, and semantically meaningful trajectories while maintaining tractability. Our project page is available at https://hta-diffusion.github.io.
I Introduction
Diffusion models have garnered increasing attention in robotic motion planning and shown strong flexibility across tasks [1, 2, 3, 4, 5]. However, most existing approaches adopt a default corruption process using zero-mean, isotropic Gaussian noise, overlooking the inherent structure of robotic trajectories. In contrast to image or text data, robotic motion possesses intrinsic properties such as temporal smoothness (e.g., consistent evolution across time steps for dynamic feasibility) and can also be shaped by structured external information, such as task-imposed conditions (e.g., start and goal states or phase-specific constraints). These structured sources of information provide inductive biases that diffusion models can exploit, most commonly through conditioning or guidance [3, 6, 7, 8, 1].
In this work, we propose to encode such structure in the form of a task-aware noise model. This design enables our diffusion planner to respect smoothness and constraint satisfaction by construction. We realize this idea with a two-level hierarchical diffusion framework. The upper level uses a denoising diffusion probabilistic model (DDPM) [9] with isotropic Gaussian noise to produce sparse, task-centric key states. For example, in navigation tasks, key states can be chosen as evenly spaced waypoints between the start and goal. In manipulation tasks, they naturally align with task phases (e.g., approach, grasp, lift), augmented by a binary contact-flag trajectory. [^1] Hereafter, we use key states to denote both waypoints in navigation and keyframe states in manipulation. We treat upper-level key states as soft observations with explicit uncertainty, so these imperfect predictions guide sampling without hard constraints or rejection. The lower-level diffusion then generates the full trajectory under a task-conditioned Gaussian, obtained by conditioning a linear time-varying Gaussian process (LTV-GP) motion prior on the key states, inspired by Gaussian process motion planning (GPMP) [10]. This yields a non-isotropic and temporally correlated noise model whose mean and covariance encourage smooth dynamics between key states, while the key states act as anchors that prevent global oversmoothing and allow the trajectory to adapt at sharp turns or task-phase changes. Accordingly, this framework produces temporally consistent and constraint-aware plans.
Prior diffusion works typically keep the corruption isotropic and inject structural constraints through conditioning or guidance during inference [8, 1, 2]. For instance, conditioning has been used to incorporate visual context from observations or to inject goal information [3, 6]. Guidance mechanisms and auxiliary objectives incorporate dynamics, task rewards, or safety specifications into the generative process [11, 12]. In addition to conditioning and guidance, task-centric representations have been leveraged to encode constraints, for example object-centric demonstrations [13] or SE(3)-based cost fields that couple grasp and motion optimization [14]. Projection-based methods have also been explored, where diffusion samples are mapped into feasible sets such as dynamically admissible trajectories [15] or collision-free multi-robot paths [16]. Beyond these categories, diffusion models have also been integrated with model-based motion planning [17] and with neural dynamics models [18].
Structured corruptions during training have been explored primarily outside robotics and remain rare within it. Examples include non-Gaussian but zero-mean choices [19], per-pixel variance schedules without temporal or robotics priors [20], and non-isotropic formulations that capture joint dependencies in human pose but still assume zero-mean noise and omit temporal correlations along the horizon, limiting their use for trajectory state conditioning [21]. Mixture priors have also been proposed for multimodality and controllability, where Gaussian mixture parameters may be predefined through clustering [22] or heuristic structure [23], learned end-to-end to adaptively cover target support [24], or dynamically adjusted during inference by solver-based approaches [25]. In robotics, structured corruptions remain rare, with only a few recent efforts—for example, integrating coarse trajectory predictions so that diffusion acts as a refiner [26], introducing temporally correlated noise such as colored Gaussian perturbations [27], or employing diffusion bridges that initiate denoising from informative prior actions in navigation tasks [28]. While these studies highlight the benefits of structured corruptions for distributional expressiveness, they do not leverage structured priors to encode task conditions and temporal smoothness simultaneously. In contrast, our design directly biases the generative process toward trajectories that are both dynamically consistent and task-aware, rather than only improving multi-modality.
Hierarchical diffusion has been used to separate coarse-level decisions from finer-level motion, for example: subgoal generation followed by refinement [29], cascaded global and local prediction with online repair [30], progressive coarse-to-fine refinement processes that achieve real-time planning frequencies [31], contact-guided manipulation policies that decompose high-level contact planning from low-level trajectory generation [32], kinematics-aware frameworks with high-level end-effector prediction and low-level joint generation [33], and diffusion–flow hybrids that pair high-level diffusion with low-level rectified flow [34]. A common pattern feeds high-level outputs to the low level as conditioning, while low-level corruption remains isotropic. This leaves the forward process and loss unchanged, so temporal coupling, boundary consistency, and waypoint fidelity must be relearned at every step. Our hierarchy instead reparameterizes the low-level prior by conditioning a GP on upper-level key states. This yields a trajectory posterior whose mean interpolates the key states and whose covariance couples time steps, so reverse updates follow a Mahalanobis geometry that narrows the search to trajectories consistent between key states and treats imperfect key states probabilistically. This improves feasibility without post hoc constraints or rejection.
We evaluate on Maze2D goal reaching and KUKA block stacking from the Diffuser benchmarks [2]. We compare (i) a single layer diffuser with isotropic noise, (ii) a hierarchical diffuser that only conditions on upper-level outputs while keeping isotropic noise, and (iii) our full hierarchical model that constructs the structured prior from upper-level key states and denoises under that prior. Across tasks, our model (iii) outperforms the conditioning-only baseline (ii), indicating that conditioning alone is insufficient. Embedding structure in the generative noise yields consistent gains in success and collision or constraint satisfaction.
Our contributions are summarized as follows:
  • A two-level hierarchical diffusion planner where the upper-level DDPM infers sparse, task-centric key states and the lower level denoises under a GPMP-based motion prior conditioned on these states. Key states are treated as soft observations with explicit uncertainty, encoding smoothness, boundary and phase consistency, and handling non-stationarities without hard constraints.
  • A diffusion formulation with task-conditioned non-isotropic Gaussian corruption obtained by conditioning a GPMP-based prior on key states. We derive closed-form forward marginals and posteriors for variational training, yielding a Mahalanobis loss that generalizes MSE while preserving the standard DDPM pipeline.
  • A conceptual and empirical comparison to conditioning-only diffusion hierarchies. Reparameterizing the low-level prior alters the diffusion landscape and concentrates sampling on trajectories consistent with key states, improving feasibility without extra constraints or loss terms.
  • Experiments on Maze2D goal reaching and KUKA block stacking showing higher success and better constraint satisfaction than baselines with isotropic noise or conditioning alone, with ablations demonstrating the benefits of the structured prior.
Figure 1: Overview of our hierarchical diffusion framework. Training (left): From dataset trajectories we extract task-relevant key states and timings $(Y,C)$. These supervise the upper-level diffusion (Alg. 1) and also define the structured prior for the lower level, trained under this task-aware noise model (Alg. 2). Test-time planning (right): For a new task specification, the upper level provides $(Y,C)$ via rule-based heuristics or the learned diffusion. These outputs instantiate the task-specific prior, from which the lower level initializes and denoises to generate a full trajectory (Alg. 3).
II Preliminary
II-A DDPM for motion planning
Let a robot trajectory be $\tau=[x_{1},\ldots,x_{H}]\in\mathbb{R}^{H\times d}$, where $x_{t}\in\mathbb{R}^{d}$ is the robot state at time step $t$ and $H$ is the planning horizon. We denote by $\tau^{i}$ the trajectory after $i$ forward diffusion steps.
Diffusion models learn to denoise trajectories that are gradually corrupted by Gaussian noise. Under the standard DDPM corruption [9], the forward process is
$\tau^{i}=\sqrt{\alpha_{i}}\,\tau^{i-1}+\sqrt{1-\alpha_{i}}\,\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,I),$ (1)
where $\alpha_{i}:=1-\beta_{i}$ for $i=1,\ldots,N$, $\{\beta_{i}\}_{i=1}^{N}\subset(0,1)$ is a scalar variance schedule, and $N$ is the total number of diffusion steps. Let $\bar{\alpha}_{i}:=\prod_{k=1}^{i}\alpha_{k}$. Then the closed-form marginal is
$\tau^{i}\mid\tau^{0}\sim\mathcal{N}\!\big(\sqrt{\bar{\alpha}_{i}}\,\tau^{0},\ (1-\bar{\alpha}_{i})I\big).$
The one-step forward posterior is also Gaussian,
$q(\tau^{i-1}\!\mid\!\tau^{i},\tau^{0})=\mathcal{N}\!\big(\tau^{i-1};\,\hat{\mu}_{i}(\tau^{i},\tau^{0}),\,\hat{\beta}_{i}I\big),$ (2)
where $\hat{\mu}_{i}$ and $\hat{\beta}_{i}$ are the posterior mean and variance coefficient at the $i^{\rm th}$ diffusion step (closed forms omitted for simplicity).
A denoising network $\mu_{\theta}(\tau^{i},i,\texttt{cond})$ is trained to predict the posterior mean with the objective
$\mathcal{L}(\theta)=\mathbb{E}_{q}\!\left[\big\|\hat{\mu}_{i}-\mu_{\theta}(\tau^{i},i,\texttt{cond})\big\|_{2}^{2}\right],$ (3)
where cond may include initial and goal states, maps, waypoints, or other task variables. At test time, sampling runs the reverse Markov chain starting from $\tau^{N}\sim\mathcal{N}(0,I)$.
In Sec. III, we generalize (1) by biasing the mean and allowing a non-isotropic covariance in the corruption process. The standard DDPM is recovered as $\mu=0$ and $\mathcal{K}=I$.
II-B Gaussian Process Motion Planning
Gaussian Process Motion Planning (GPMP) models a continuous time trajectory as a Gaussian process (GP), i.e.,
$\tau(t)\sim\mathcal{GP}(\tilde{\mu}(t),\tilde{\mathcal{K}}(t,t^{\prime})),$ (4)
whose mean and kernel are induced by a linear time varying (LTV) dynamics prior [10]. Consider the stochastic LTV model $\dot{x}(t)=A(t)x(t)+u(t)+w(t)$ where $w(t)$ is zero mean white noise with spectral density $Q_{C}\succeq 0$. Let the initial mean and covariance at time $t_{0}$ be $\mu_{0}$ and $\mathcal{K}_{0}$, and let $\Phi(t,s)$ denote the state transition matrix of $A(\cdot)$. Then the GP has mean $\tilde{\mu}(t)=\Phi(t,t_{0})\mu_{0}+\int_{t_{0}}^{t}\Phi(t,s)u(s)\,ds$ and kernel $\tilde{\mathcal{K}}(t,t^{\prime})=\Phi(t,t_{0})\mathcal{K}_{0}\Phi(t^{\prime},t_{0})^{\top}+\int_{t_{0}}^{\min\{t,t^{\prime}\}}\Phi(t,s)Q_{C}\Phi(t^{\prime},s)^{\top}ds$, which admit sparse and exactly computable discrete counterparts.
Following [1], our work uses a discrete-time version of (4) derived from a constant velocity or constant acceleration model with step $\Delta t$. Writing the state at timestep $t$ as $x_{t}$, the one-step evolution is $x_{t+1}=\Phi_{t,t+1}x_{t}+q_{t}$ with $q_{t}\sim\mathcal{N}(0,Q_{t,t+1})$, where $\Phi_{t,t+1}=\begin{bmatrix}I&\Delta t\,I\\ 0&I\end{bmatrix}$ and $Q_{t,t+1}=\begin{bmatrix}\tfrac{1}{3}\Delta t^{3}Q_{C}&\tfrac{1}{2}\Delta t^{2}Q_{C}\\ \tfrac{1}{2}\Delta t^{2}Q_{C}&\Delta t\,Q_{C}\end{bmatrix}$. The same structure applies, no matter if the state includes positions only or positions and velocities; velocities can be included explicitly or marginalized to obtain an equivalent GP prior over positions with a block tridiagonal precision matrix. In our experiments, Maze2D uses positions and velocities, whereas the manipulation arm uses joint positions only.
Figure 2: Trajectory generation pipeline. Input: task specification (e.g., Maze2D goal reaching; KUKA grasp-and-place) and initial conditions. Upper level produces key states (stack of which is denoted with $Y$) and their timings (equivalent to selection matrix $C$), either by diffusion module and/or rules. For Maze2D, they are waypoints with fixed timings. For KUKA, they are grasp/release states and contact flags, which implicitly imply timing. With $(Y,C)$, we construct a Gaussian prior by conditioning a GP over trajectories, yielding $(\mu,\mathcal{K})$ via (9)–(10). The shaded regions illustrate covariance structure $\mathcal{K}$ around the trajectory, showing tighter variance near key states and looser variance elsewhere. The lower level then generates the full trajectory by reverse diffusion from $\mathcal{N}(\mu,\mathcal{K})$.
III Method
III-A Diffusion with a task conditioned Gaussian and hierarchical planner overview
Let $\mathcal{D}$ be a dataset of robot trajectories where each clean sample $\tau\!\in\!\mathcal{D}$ is paired with task information that may admit multiple feasible solutions. In our settings, this task information includes start and goal or a waypoint template in Maze2D, and grasp and place timing with key joint configurations in KUKA manipulation block stacking.
It is desirable for the corruption (forward) process to retain task information so that reverse diffusion starts from a task-relevant noise prior. To this end, we modify the standard DDPM to use a non-isotropic Gaussian with a task conditioned mean and covariance. Reusing the schedule notation from Sec. II-A, with $\alpha_{i}=1-\beta_{i}$ and $\bar{\alpha}_{i}=\prod_{k=1}^{i}\alpha_{k}$, we define
$\tau^{i}\sim\mathcal{N}\!\Big(\sqrt{\alpha_{i}}\,\tau^{i-1}+(1-\sqrt{\alpha_{i}})\,\mu,\;(1-\alpha_{i})\,\mathcal{K}\Big),$ (5)
where $\mu\!\in\!\mathbb{R}^{H\times d}$ and $\mathcal{K}\!\in\!\mathbb{R}^{Hd\times Hd}$ with $\mathcal{K}\succeq 0$ encode task-dependent trajectory structure. $\tau^{0}$ is just $\tau$.
As in DDPM, (5) yields closed form conditionals with respect to $\tau^{0}$:
$\tau^{i}\mid\tau^{0}\sim\mathcal{N}\!\Big(\sqrt{\bar{\alpha}_{i}}\,\tau^{0}+(1-\sqrt{\bar{\alpha}_{i}})\,\mu,\;(1-\bar{\alpha}_{i})\,\mathcal{K}\Big).$ (6)
The one step forward posterior is also Gaussian,
$q(\tau^{i-1}\!\mid\!\tau^{i},\tau^{0})=\mathcal{N}\!\big(\tau^{i-1};\,\tilde{\mu}_{i}(\tau^{i},\tau^{0},\mu),\,\tilde{\beta}_{i}\,\mathcal{K}\big),$ (7)
with coefficients $\tilde{\mu}_{i}$ and $\tilde{\beta}_{i}$ that directly generalize the DDPM posterior mean $\hat{\mu}_{i}$ and variance coefficient $\hat{\beta}_{i}$ from Sec. II-A (explicit derivations are given in the Appendix). As $N\to\infty$, the terminal distribution approaches $\tau^{N}\!\sim\!\mathcal{N}(\mu,\mathcal{K})$, so at test time we initialize the reverse diffusion from this task-specific Gaussian.
At test time, we aim to generate a trajectory $\tau$ that is both dynamically feasible and aligned with task constraints. Consequently, the mean $\mu$ and covariance $\mathcal{K}$ of the non-isotropic Gaussian should encode both dynamics and task information. We achieve this with a two-level hierarchy. The upper level extracts task information that is available without access to the full trajectory and uses it to instantiate $(\mu,\mathcal{K})$. Concretely, it generates sparse task-centric key states with associated uncertainties that we treat as soft observations. The lower level integrates these key states with the GPMP prior (4) from Sec. II-B. Conditioning the GP prior on the soft observations yields a posterior mean and covariance that define $\mu$ and $\mathcal{K}$ in (5). During test time, reverse diffusion then proceeds under this non-isotropic and time correlated Gaussian, starting from $\mathcal{N}(\mu,\mathcal{K})$.
The overall planner therefore has two functions: the upper level produces key states and their associated timings, and the lower level generates the full trajectory by reverse diffusion given $(\mu,\mathcal{K})$. Inputs are a task specification and an initial condition. An overview of the framework is shown in Fig. 1. The next subsections detail each level.
III-B Upper level: key states and timings
The upper level identifies task-relevant key states and their timings, using a lightweight diffusion module parameterized by $\phi$, combined with simple rule-based heuristics as needed. By key states, we mean boundary conditions, waypoints, or states at contact events.
Given a full trajectory $\tau$, let $Y=[y_{1};\ldots;y_{n}]\in\mathbb{R}^{nd}$ be the stack of $n$ key states. Define a binary selection matrix $C\in\{0,1\}^{nd\times Hd}$ that extracts the key states from the full trajectory, so that $Y=C\,\tau$. Note that the selection matrix $C$ is equivalent to the timings for the key states $Y$. The upper level generates these values, and the construction of them depend on the task, as described below.
Maze2D. Key states are evenly spaced temporal waypoints between the start and the goal, so $C$ is fixed by design. In this example, the upper-level diffusion is trained to generate the waypoint set $Y$.
KUKA manipulation. Key states are defined as the joint configuration at the grasp and release timings. The upper-level diffusion is trained to predict the grasp and release timings by generating a binary contact-flag trajectory (1 for contact, 0 for no contact). From this trajectory, we can obtain $C$. The corresponding joint-state set $Y$ is constructed using simple rules (e.g., block-selection heuristics, detailed in Sec. IV) together with inverse kinematics (IK).
The lower level is agnostic to how the high level selects key states. Even if the high level produces imprecise or partially inconsistent $(Y,C)$ with the full trajectories, the lower level treats them as soft observations with uncertainty and proceeds unchanged. Through GP conditioning, this uncertainty enters as observation covariance, so imperfect inputs are down-weighted rather than enforced as hard constraints, and the reverse diffusion remains stable and feasible. For more details, please refer to the subsection below.
III-C Lower level: trajectory generation by denoising
Given the key states $Y\in\mathbb{R}^{nd}$ and the corresponding selection matrix $C\in\{0,1\}^{nd\times Hd}$, we treat $Y$ as soft observations of the stacked trajectory $\tau=[x_{1};\ldots;x_{H}]\in\mathbb{R}^{H\times d}$ with observation covariance $\mathcal{K}_{y}\succeq 0$. The unconditioned GPMP prior (4) induces a GP prior over $\tau$. Conditioning this GP prior to the soft observations $(Y,C,\mathcal{K}_{y})$ yields a task conditioned posterior over the full trajectory,
$\tau(t)\sim\mathcal{N}\big(\mu(t),\,\mathcal{K}(t,t^{\prime})\big),$ (8)
with a closed form mean and covariance [10]
[…]
$\mathcal{K}$ $=\widetilde{\mathcal{K}}-\widetilde{\mathcal{K}}\,C^{\top}\!\left(C\,\widetilde{\mathcal{K}}\,C^{\top}+\mathcal{K}_{y}\right)^{-1}\!C\,\widetilde{\mathcal{K}}$. (10)
Small variances in $\mathcal{K}_{y}$ enforce tight adherence to key states (e.g., fixed initial or goal states), while larger variances allow flexibility (e.g., soft waypoints). Intuitively, $Y$ anchors the trajectory near important states and $\mathcal{K}_{y}$ encodes confidence in those anchors, without imposing hard constraints.
The conditioned GP (8) determines the non-isotropic and time-correlated Gaussian noise used in our diffusion model; at each step, the injected noise is drawn with a mean and covariance shaped by linear time varying dynamics and the key state and timing information. In particular, the lower level initializes reverse diffusion from $\mathcal{N}(\mu,\mathcal{K})$ and denoises under the general forward and posterior relations introduced in Sec. II-A. Recall that the one-step forward posterior for our biased and non-isotropic setting has the closed form
$q(\tau^{i-1}\!\mid\!\tau^{i},\tau^{0})=\mathcal{N}\!\big(\tau^{i-1};\,\tilde{\mu}_{i}(\tau^{i},\tau^{0},\mu),\,\tilde{\beta}_{i}\,\mathcal{K}\big),$ (11)
where $\tilde{\mu}_{i}$ and $\tilde{\beta}_{i}$ generalize the DDPM posterior coefficients $\hat{\mu}_{i}$ and $\hat{\beta}_{i}$. For learning we approximate the reverse kernel
$p_{\theta}(\tau^{i-1}\!\mid\!\tau^{i})=\mathcal{N}\!\big(\tau^{i-1};\,\mu_{\theta}(\tau^{i},i,\texttt{cond}),\,\Sigma^{i}\big),$
where cond may include $Y$, $C$, initial states, and other context, and we take $\Sigma^{i}$ to match the scaled structure $\tilde{\beta}_{i}\,\mathcal{K}$. Following [9] and Sec. II-A, minimizing the KL divergence between $p_{\theta}$ and the true forward posterior reduces to a weighted regression toward $\tilde{\mu}_{i}$. Using the Mahalanobis norm $\|v\|^{2}_{A}:=v^{\top}A\,v$, the training objective becomes
$L(\theta)=\mathbb{E}_{q}\!\left[\big\|\tilde{\mu}_{i}-\mu_{\theta}(\tau^{i},i,\texttt{cond})\big\|^{2}_{\mathcal{K}^{-1}}\right],$ (12)
up to a scalar schedule dependent factor. Compared to the isotropic MSE, (12) naturally weights errors by the inverse of the task conditioned covariance, encouraging denoising steps that stay consistent with the GP induced temporal coupling and the confidence encoded in $\mathcal{K}_{y}$.
III-D Pipeline Summary
The complete training process is shown in Alg. 1 and Alg. 2. For the upper level, we first obtain the ground truth $(C,Y)$ and choose the supervision targets, which are task dependent We then update the network parameters of the upper-level diffusion with the standard diffusion loss on the chosen target(s) (Alg. 1). For the lower level, we use the ground truth $(C,Y)$ to instantiate the task specific prior by computing $(\mu,\mathcal{K})$ via (9)–(10). We then update the network parameters of the lower-level diffusion with the Mahalanobis MSE (Alg. 2).
The complete inference process is shown in Alg. 3. At sampling time, the upper level outputs $(C,Y)$, from which we construct $(\mu,\mathcal{K})$. At the lower level, we use $(\mu,\mathcal{K})$, initialize $\tau^{N}\sim\mathcal{N}(\mu,\mathcal{K})$, and run the reverse chain using (7) with the learned denoiser until a fully denoized trajectory is obtained.
Remark 1
Computing covariance-related terms $\big(C\widetilde{\mathcal{K}}C^{\top}+\mathcal{K}_{y}\big)^{-1}$, $\mathcal{K}$, and $\mathcal{K}^{-1}$ repeatedly can be costly. We therefore precompute the corresponding gain terms for all timing configurations used at training and inference. For example, in the KUKA manipulation task, we divide the horizon $H$ into 10 bins and precompute gains for all grasp and release bin pairs $\tbinom{10}{2}=45$, since only $C$ changes with timing choices. This preserves time range conditioning while keeping online costs low.
Algorithm 1 Upper-level training (for key states and/or timings)
1:Inputs: dataset $\mathcal{D}$ of trajectories $\tau$
2:Initialize upper-level diffusion module parameters $\phi$
3:while not converged do
4:  Sample $\tau\sim\mathcal{D}$
5:  Extract ground-truth supervision targets $(C,Y)$ $\triangleright$ depending on application, train on $C$, $Y$, or both
6:  Update $\phi$ using the standard diffusion loss w.r.t. the chosen target(s)
7:end while
Algorithm 2 Lower-level training (for full trajectory)
[…]
2:Initialize lower-level diffusion module parameters $\theta$
3:while not converged do
4:  Sample $\tau\sim\mathcal{D}$, set $\tau^{0}=\tau$, and extract ground truth $(C,Y)$
5:  Form $(\mu,\mathcal{K})$ via (9)–(10) (usually extract from cached pre-computed ones)
6:  Sample $i\sim\text{Unif}\{1,\dots,N\}$
7:  $\tau^{i}\!\sim\!\mathcal{N}\!\left(\sqrt{\bar{\alpha}_{i}}\tau^{0}+(1-\sqrt{\bar{\alpha}_{i}})\mu,\;(1-\bar{\alpha}_{i})\mathcal{K}\right)$
8:  Gradient descent on $\nabla_{\theta}\left\|\tilde{\mu}_{i}(\tau^{i},\tau^{0})-\mu_{\theta}(\tau^{i},i)\right\|_{\mathcal{K}^{-1}}$
9:end while
Algorithm 3 Test-time Planning: Hierarchical Diffusion
[…]
2:Upper level: produce $(C,Y)$ $\triangleright$ from diffusion model $\phi$ and rule-based heuristic
3:Prior build: $(\mu,\mathcal{K})\leftarrow\text{GP-Post}(\tilde{\mu},\widetilde{\mathcal{K}},C,Y,\mathcal{K}_{y})$ via (9)–(10)
[…]
5:for $i=N,\dots,1$ do
6:  $\tau^{i-1}\sim\mathcal{N}(\mu_{\theta}(\tau^{i},i),\tilde{\beta}_{i}\mathcal{K})$
7:end for
8:return $\tau^{0}$
Remark 2
Motion Planning Diffusion (MPD) [1] applies a GP as a guidance cost only during reverse sampling, while the forward corruption remains isotropic. We instead parameterize the noise with a GP: conditioning on key states sets the mean $\mu$ and covariance $\mathcal{K}$ that govern both the forward and posterior distributions. This yields three benefits. First, structure is present at every step of training and sampling, so denoising follows a Mahalanobis geometry that narrows the search to trajectories consistent with temporal coupling and task constraints. Second, uncertainty in key states enters as an observation covariance $\mathcal{K}_{y}$, which downweights imperfect cues without hard constraints. Third, feasibility comes from the generative model rather than from external costs, reducing sensitivity to guidance weights and improving stability over long horizons.
Figure 3: Reverse diffusion in Maze2D. Columns show intermediate denoising steps (left to right). Top: isotropic Gaussian prior—samples start fully random and remain noisy until late steps. Bottom: our model—initialization is already task-aware and trajectories quickly converge to smooth, goal-reaching solutions.
Figure 4: Maze2D goal reaching. Top-Left: task map with start (blue) and goal (red). Top-right: baseline (isotropic Gaussian noise) samples. Bottom: our method. The small inset on the left shows the waypoints predicted by the upper level (see Fig. 2). Although waypoints provide a rough guide for the global trajectory, they can be imperfect or even infeasible (with collisions), so hard-constraining the lower level using these waypoints is brittle. Thus, we treat them as noisy observations and encode their uncertainty in the GPMP prior with the observation covariance $\mathcal{K}_{y}$, which guides denoising to yield smooth, collision-free trajectories aligned with the task.
Figure 5: Comparison of generated trajectories for KUKA block stacking. Each row shows a task with a random initial arm configuration and block positions. Columns: (left) initial condition, (middle) trajectories generated with an isotropic Gaussian prior, (right) trajectories generated with our model.
IV Experiments
We adopt Diffuser [2] as the baseline and evaluate the proposed structured-noise formulation on two domains: (i) Two-dimensional maze navigation (Maze2D) with initial and goal conditioning, and (ii) KUKA block stacking with grasp and release events represented with contact flag trajectory for each block. We report goal success, dynamic feasibility, and constraint satisfaction.
IV-A Experimental Setup
IV-A1 Comparison models
We compare four planners that share the same network backbone and differ only in how they use key states and in the structure of the diffusion noise:
  • Isotropic (without key states). A standard DDPM with isotropic Gaussian corruption and no high-level key states (the flat Diffuser baseline).
  • Isotropic (with key states as conditioning). A standard DDPM at the lower level that receives the high-level key states as conditioning inputs, but still uses isotropic Gaussian corruption at the low level.
  • GPMP prior (without key states). A single-level diffuser with non-isotropic Gaussian noise shaped by the GPMP prior, but without key states. This highlights that relying only on a GP prior for dynamic feasibility can harm trajectory generation by causing oversmoothing.
  • GPMP prior (with key states). Our method: a two-level hierarchy in which the upper level produces task-centric key states and timings, and the lower level denoises under the non-isotropic Gaussian from GPMP.
Note that the choice of noise prior applies only to the lower-level diffusion. Whenever a hierarchy is used, the upper-level diffusion remains a standard DDPM with isotropic Gaussian noise.
IV-A2 Neural network architecture
Both levels use a UNet architecture as in [2]. For the lower level, the network is conditioned on the key states and their timings produced by the upper level. Apart from this conditioning and the choice of noise model (isotropic versus GP-conditioned), the architecture and training schedule are identical across all models so that any performance differences are induced by the effect of the noise structure.
IV-A3 Maze2D goal reaching
Each task specifies a start and a goal state on a fixed maze map. The upper-level diffusion produces temporal waypoints $Y$ at fixed times $C$, which are treated as soft observations with associated uncertainty. These waypoints instantiate the task-conditioned GP prior, from which the lower level generates full trajectories by reverse diffusion.
IV-A4 KUKA arm block stacking
The task is grasp-and-place with a 7-DoF KUKA arm. The dataset provides 7-DOF joint position trajectories and a binary contact-flag trajectory (one bit per time step) for each block, while block poses are determined deterministically from kinematics and contact.
As explained in Sec.III, the key states are grasp and release joint positions. Given a ground-truth trajectory $\tau^{0}$ and its contact flags, we define the grasp configuration as the joint state at the first $0\!\to\!1$ transition and the release configuration as the joint state at the first $1\!\to\!0$ transition. The stack $Y$ consists of the initial joint state together with these grasp and release joint states, while $C$ is taken to be the binary contact-flag trajectory over the horizon. The upper-level diffusion is trained to generate $C$ conditioned on $Y$.
At test time, we input the grasp and stack targets with the following heuristics: (1) if a block is currently grasped, use it as the grasp target; (2) choose the release target as the block with the largest vertical position, excluding any grasped block; (3) if no block is grasped, choose the block closest to the initial end-effector position; (4) if the chosen grasp–release pair is already stacked ($|dx|<\mathrm{tol}$, $|dy|<\mathrm{tol}$), select a different grasp block. After the block selection, we compute the grasp and release joint states by IK to form $Y$. Then, the upper-level diffusion predicts the contact-flag trajectory $C$; grasp and release timings are extracted from the first $0\!\to\!1$ and $1\!\to\!0$ transitions, respectively.
IV-B Results
IV-B1 Success Rate
In Maze2D, we define a trajectory as successful if two conditions are met: (i) the final state lies within a predefined tolerance radius around the goal, and (ii) the entire trajectory remains collision-free with respect to the maze map. For each planning model, results are averaged over 100 randomly sampled start–goal pairs.
In KUKA block stacking, success requires both a valid grasp and a valid placement within tolerance. We declare a grasp successful when the end-effector is within a spatial threshold of the target block at the first contact onset (flag transition $0\!\to\!1$). A placement is declared successful when, at the release time (flag transition $1\!\to\!0$), the end-effector is within tolerance of the designated stack location. Results are averaged over 100 randomized initial joint configurations and block layouts.
We evaluate a $2\times 2$ grid of diffusion planners that vary along two axes: (i) isotropic Gaussian noise versus the GPMP prior (structured, task-aware mean $\mu$ and covariance $\mathcal{K}$), and (ii) without versus with high-level key states. All models share the same UNet backbone, are trained for the same number of optimization steps, and are verified to have comparable convergence in training loss.
Tables I and II report the success rates. Across both tasks, the best performance is achieved by combining the upper-level key-state generator with the GPMP prior. Intuitively, upper-level key states and timings supply task semantics (e.g., guidance through turns in a maze or contact phases in manipulation), while the GPMP prior encodes smoothness and soft constraint satisfaction. Together, these two components concentrate probability mass of the reverse process near dynamically feasible and task-consistent manifolds, thereby reducing the burden on the denoiser.
Interestingly, the GPMP prior without informative key states underperforms even the isotropic baseline. This occurs because, although the GPMP prior captures temporal correlations and promotes smoothness, it can oversmooth when applied uniformly across the entire horizon. Realistic trajectories often involve non-stationarities: sharp turns in navigation and abrupt grasp/release events in manipulation. Without explicit key states to anchor the dynamics, the GP prior may bias trajectories away from the task-critical states, leading to degraded success. Effective planning, therefore, requires a balance: the hierarchical key-state generator anchors task-critical events, while the prior provides tighter variance around the key states ensuring they are roughly respected, while maintaining smoothness and temporal consistency. The combination of both yields robust, dynamically feasible, and task-aligned plans.
Maze2D (100 samples) Success
Isotropic (without key states) 36/100
Isotropic (with key states as conditioning) 42/100
GPMP Prior (without key states) 14/100
GPMP Prior (with key states) (our model) 75/100
TABLE I: Maze2D success rates (number of collision-free successful trajectories out of 100) after 8080k training steps. Model names match the definitions in Sec. IV.
KUKA stacking (100 samples) Success
Isotropic (without key states) 11/100
Isotropic (with key states as conditioning) 12/100
GPMP Prior (without key states) 8/100
GPMP Prior (with key states) (our model) 71/100
TABLE II: KUKA stacking success rates (number of successful grasps and placements out of 100) after 3030k training steps. Model names match the definitions in Sec. IV.
Figure 6: Maze2D velocity consistency. Left: isotropic Gaussian prior; Right: our model. Blue: generated velocities; Orange: numerical derivatives of generated positions. Closer alignment indicates better dynamic feasibility, achieved with the GPMP prior.
Maze2D (100 samples) Mean Absolute Error
Isotropic (without key states) 0.22
Isotropic (with key states as conditioning) 0.21
GPMP Prior (without key states) 0.08
GPMP Prior (with key states) (our model) 0.08
TABLE III: Maze2D velocity consistency: mean absolute error (MAE) between generated velocities and finite-difference derivatives of positions over 100 trajectories. With GPMP noise, MAE remains low regardless of key states, since the LTV-GP prior enforces natural position–velocity alignment.
IV-B2 Dynamic Feasibility & Smoothness
We next evaluate whether the generated trajectories satisfy dynamic feasibility and exhibit smoothness.
Maze2D.
The Diffuser baseline provides position, velocity, and action effort sequences. While the success rate is computed only from position plans, our method is trained with a GPMP-inspired prior that couples position and velocity, thereby encoding feasibility. To assess this property, we compare the generated velocities against numerical derivatives of the generated positions (finite difference over the time step, Fig. 6). We also report the mean absolute error (MAE) along the horizon (Table III). We additionally present representative trajectories (Fig. 4) along with their denoising processes (Fig.  3), providing qualitative insight into the behavior of the models. Qualitatively, our method generates smooth and feasible motions, whereas the isotropic baseline often produces jagged position traces with discontinuities inside the trajectories (caused by hard enforcement of start and goal conditions), resulting in unrealistic and infeasible plans .
KUKA block stacking.
The open-source baseline does not provide velocities or actions. Thus, feasibility cannot be directly evaluated in this domain, but smoothness can still be assessed by measuring the proximity between consecutive joint states along the trajectory. Visualizations confirm that trajectories generated with our structured prior are smoother and more coherent compared to the baseline (Fig. 5).
V Discussion and Conclusion
Embedding GPMP structure in the noise yields temporally consistent, constraint-aware trajectories, improving success under sparse or imperfect high-level guidance. Our framework is task-centric: the upper level extracts task-relevant key states/timings, and the lower-level diffusion uses the GP-conditioned $(\mu,\mathcal{K})$ to bias sampling toward feasible solutions. We validate the approach through experiments on two robotic domains. In Maze2D, the prior concentrates probability mass near start/goal and along smooth waypoint-guided paths; in KUKA stacking, predicted contact timing shapes the kernel so denoising honors grasp/release phases. Although we proposed a lightweight diffusion model for the upper level, the lower level is agnostic to how the key states are obtained—be it rule-based, model-based, or learned. It always treats $(Y,C)$ as soft observations with uncertainty; GP conditioning down-weights imperfect inputs rather than enforcing them as hard constraints, leaving room to pair this framework with alternative task specifiers or high-level motion planners.
Moreover, our method does not require handcrafted reward terms or auxiliary objectives, and no advanced neural architectures are used beyond conditioning on upper-level-generated features. Thus improvements in sample quality and task success arise from where the structure lives—in the noise—rather than from heavier networks or guidance heuristics. This shows our method’s efficacy even with simple mechanics.